# Is Compression Really Linear with Code Intelligence?

### v2

@@ 2.2 Code Intelligence
The remarkable capabilities of LLMs in code-related tasks have spurred the concept of “code intelligence”—the comprehensive ability to understand, generate, and manipulate code (rabin2021understanding, ; wan2024deep, ). Early evaluations typically focused on Python code completion tasks, such as HumanEval and MBPP (zhu-etal-2024-survey, ). However, as code applications and workflows grow in complexity (wang2024agents, ; roberttransforming, ), it is increasingly evident that a robust notion of code intelligence must extend beyond unilingual assessments to encompass proficiency across multiple programming languages (cassano2022multipl, ; peng-etal-2024-humaneval, ; chai2024mcevalmassivelymultilingualcode, ; liu2024mdeval, ). Furthermore, a growing body of research suggests that comprehensive code intelligence is not a monolithic entity but consists of a set of fundamental, or “atomic”, tasks. While existing taxonomies for code-related LLM tasks vary—for instance, by input/output modalities (e.g., NL-PL, PL-NL) or by user intent (chen2024survey, ) —they converge on a core set of underlying capabilities. We identify these as: (1) Code Generation: automated creation of source code from natural language descriptions (chen2021evaluating, ; austin2021program, ; zhuo2024bigcodebench, ). (2) Code Explanation (or Understanding): grasp and articulate the syntax, semantics, functionality, and underlying intent of a given code snippet (meyerson2025position, ; fang2024large, ; nam2024using, ). (3) Code Reasoning: infer code properties and predict code behavior without execution (dehghan2024assessing, ; chen2024reasoning, ). (4) Code Repair: identify and generate patches to correct bugs, vulnerabilities, or other errors in source code (pasquale2025challenges, ; liu2024mdeval, ). While “code intelligence” is a broad concept, its decomposition into these four atomic tasks provides a more structured and actionable framework for quantitative evaluation.
@@ 4.1 Setup
Our format annealing training set consists of 4,995 QA pairs. We use DeepSpeed ZeRO-3 optimization with BF16 precision on 8*A100-80GB GPU. The batch size, learning rate li2025predictablescalei , and warm-up ratio are set to 128, 1e-4, and 0.1, respectively. Each models are trained for up to 4 epochs. After each epoch, the ratio of successful extraction from outputs is evaluated. When it exceeds 99%, the training is stopped and the current checkpoint is used for the OmniCodeBench evaluation.
@@ 6 Conclusion
Appendix
A. Data Synthesis Pipeline for Train Set
The training set integrates three primary sources: CommitPackFT [41], CodeSearchNet [27], and LeetCode, each subjected to a specialized preprocessing workflow. The CommitPackFT data containing paired “old contents” and “new contents” code revisions serves as the foundation for the code fix task. From an initial pool of approximately 50,000 sampled instances, we implemented a multistage cleaning protocol: first removing entries where old and new contents are identical or contain empty inputs, followed by filtering out files with non-executable placeholders (e.g., “TODO”, “FIXME”). We subsequently leveraged GPT-4o to automatically verify that the code in old content reliably triggers run-time errors while their corresponding new content resolve these issues, preserving associated error logs for prompt generation. A final manual inspection phase eliminates ambiguous or meaningless cases, resulting in 3,855 rigorously validated samples. These curated instances are further standardized using GPT-4o to match the input-output formatting specifications of LiveCodeBench code fix benchmark, ensuring seamless integration with downstream evaluation frameworks.
The LeetCode dataset comprises 3,077 programming challenges and their corresponding solutions scraped from LeetCode international platform, exhibiting inherently high data quality due to its curated nature. To clean the data, we implemented a standardized filtering protocol: (1) excluding video-based solutions; (2) randomly selecting a solution for each problem; (3) structurally aligning each instance with LiveCodeBench’s code generation task format with the assistance of GPT-4o. The standardized prompt template integrates four key components: the problem statement, illustrative examples (including input and output cases), algorithmic constraints and the initial function signature, thereby preserving the essential context required for reproducible code generation tasks.
The CodeSearchNet dataset extracts function docstrings as instructions, with a high degree of structural repetition. We removed entries where the output character length exceeds 5,000, as well as those with too short or meaningless instructions. The data was then sorted in descending order by length, and 5-gram deduplication was applied to removing any instructions or output containing “TODO” or “FIXME”. We used GPT-4o to filter suitable data for each language, and finally, a comprehensive sampling was performed to obtain 8,699 data samples. Ultimately, we sampled 4,995 training data samples from three data sources as our training set.
B. Why Choose LogC
Figure 7: Linear relationship between BPC and code intelligence on our open-source validation set.
Figure 7 shows that with a RMSE of 0.115 the linear fitting curve performs worse than the logarithmic relationship curve with a RMSE of only 0.057. Therefore, in our paper the values of code intelligence are all pre-processed in logarithmic scale.
C. Individual Fitting Curve of Each Benchmark
The fitting results of each benchmark are displayed in Figure 8,  9,  10,  11,  12,  13,  14,  15,  16,  17 which seperately represent HumanEval, MBPP, McEval, BigCodeBench-Full, BIRD, Spider, LiveCodeBench-Fix, CRUXEval-I, CRUXEval-O and HumanEvalExplain.
Figure 8: Fitting curve between HumanEval benchmark and its corresponding metric.
Figure 9: Fitting curve between MBPP benchmark and its corresponding metric.
Figure 10: Fitting curve between McEval benchmark and its corresponding metric.
Figure 11: Fitting curve between BigCodeBench-Full benchmark and its corresponding metric.
Figure 12: Fitting curve between BIRD benchmark and its corresponding metric.
Figure 13: Fitting curve between Spider benchmark and its corresponding metric.
Figure 14: Fitting curve between LiveCodeBench-Fix benchmark and its corresponding metric.
Figure 15: Fitting curve between CRUXEval-I benchmark and its corresponding metric.
Figure 16: Fitting curve between CRUXEval-O benchmark and its corresponding metric.
Figure 17: Fitting curve between HumanEvalExplain benchmark and its corresponding metric.
D. Details of Open-Source Validation Set
Detailed data quantity and proportion of our open-source validation set are shown in Table 2.
Table 2: More details of our open-source validation set.
Language Number of Data Entries Number of Characters Proportion of Characters
Python 1, 500 7, 086, 432 41.21%
Go 500 1, 598, 829 9.30%
R 300 1, 460, 408 8.49%
C 100 714, 577 4.16%
C++ 200 861, 125 5.01%
Java 500 2, 020, 641 11.75%
JavaScript 500 1, 866, 524 10.85%
HTML 100 892, 826 5.19%
SQL 200 695, 673 4.05%
Total 3, 900 17, 197, 035 100.00%

### v4

@@ 2.2 Code Intelligence
The remarkable capabilities of LLMs in code-related tasks have spurred the concept of “code intelligence”—the comprehensive ability to understand, generate, and manipulate code (rabin2021understanding, ; wan2024deep, ). Early evaluations typically focused on Python code completion tasks, such as HumanEval and MBPP (zhu-etal-2024-survey, ). However, as code applications and workflows grow in complexity (wang2024agents, ; roberttransforming, ), it is increasingly evident that a robust notion of code intelligence must extend beyond unilingual assessments to encompass proficiency across multiple programming languages (cassano2022multipl, ; peng-etal-2024-humaneval, ; chai2024mcevalmassivelymultilingualcode, ; liu2024mdeval, ). Furthermore, a growing body of research suggests that comprehensive code intelligence is not a monolithic entity but consists of a set of fundamental, or “atomic”, tasks. While existing taxonomies for code-related LLM tasks vary—for instance, by input/output modalities (e.g., NL-PL, PL-NL) or by user intent (chen2024survey, ) —they converge on a core set of underlying capabilities. We identify these as: (1) Code Generation: automated creation of source code from natural language descriptions (chen2021evaluating, ; austin2021program, ; zhuo2024bigcodebench, ). (2) Code Explanation (or Understanding): grasp and articulate the syntax, semantics, functionality, and underlying intent of a given code snippet (meyerson2025position, ; fang2024large, ; nam2024using, ). (3) Code Reasoning: infer code properties and predict code behavior without execution (dehghan2024assessing, ; chen2024reasoning, ; li2025system, ). (4) Code Repair: identify and generate patches to correct bugs, vulnerabilities, or other errors in source code (pasquale2025challenges, ; liu2024mdeval, ). While “code intelligence” is a broad concept, its decomposition into these four atomic tasks provides a more structured and actionable framework for quantitative evaluation.
@@ 4.1 Setup
Our format annealing training set consists of 4,995 QA pairs. We use DeepSpeed ZeRO-3 optimization with BF16 precision on 8*A100-80GB GPU. The batch size, learning rate, and warm-up ratio are set to 128, 1e-4, and 0.1, respectively. Each models are trained for up to 4 epochs. After each epoch, the ratio of successful extraction from outputs is evaluated. When it exceeds 99%, the training is stopped and the current checkpoint is used for the OmniCodeBench evaluation.